### v3

@@ Abstract
Tidal energy is one of the key components in increasing the penetration of renewable energy. High tidal energy penetration into the electrical grid depends on accurate tidal current speed forecasting. Model inaccuracies hinder forecast accuracy. Previous research primarily used physical models to forecast tidal current speed, yet tidal current variations influenced by the orbital periods of celestial bodies make accurate physical modeling challenging. Research on the multi-periodicity of tides is crucial for forecasting tidal current speed. We propose the Wavelet-Enhanced Convolutional Network to learn multi-periodicity. The framework embeds intra-period and inter-period variations of one-dimensional tidal current data into the rows and columns, respectively, of a two-dimensional tensor. Then, the two-dimensional variations of the sequence can be processed by convolutional kernels. We integrate a time-frequency analysis method into the framework to further address local periodic features. Additionally, to enhance the framework’s stability, we optimize the framework’s hyperparameters with the Tree-structured Parzen Estimator. The proposed framework captures multi-periodic dependencies in tidal current data. Numerical results show a 10-step average Mean Absolute Error of 0.025, with at least a 1.18% error reduction compared to other baselines. Further ablation studies show a 1.4% reduction in Mean Absolute Percentage Error on the data with artificially added periodic fluctuations.
@@ 1.1 Background
The development of renewable energy can promote emission reduction [1]. Tidal energy is considered one of the most competitive forms of renewable energy [2].
Tidal energy attracts the attention of many coastal countries [3], especially the UK. The potential of tidal energy lies in three aspects: cost-effectiveness, stability, and predictability. As tidal energy gains more attention, its cost-effectiveness is expected to improve, as shown in Figure 1. Tidal energy may replace other renewable energies when the Levelized Cost of Energy (LCOE) falls below £49-55/MWh (expected by 2050) [4]. Additionally, tidal energy offers high stability, which is crucial for the economic feasibility of any renewable energy project [5]. Therefore, the introduction of tidal energy can enhance grid stability. Moreover, tidal energy provides the advantage of predictability over large time scales. Consequently, tidal energy is an excellent choice as a base-load energy supply [2].
Tidal current turbines generate electricity using the kinetic energy of tidal currents, like wind turbines use the kinetic energy of air [6]. Therefore, fluctuations in Tidal Current Speed (TCS) reduce grid stability [7]. Enhancing the accuracy of short-term forecasting helps mitigate the impact of TCS fluctuations on grid stability, thereby increasing the penetration of tidal energy into the grid [8]. Research [9] proposes to explore the statistical properties of real-life time series and design customized deep learning models for high-performance time series forecasting. The research on statistical properties of time series inspires us to explore the multi-periodicity within the TCS series. We increase the consideration of the statistical properties of TCS to improve the accuracy of TCSF.
@@
Nomenclature
$\beta$ Output length $\boldsymbol{\omega}$ A hyperparameter combination $\boldsymbol{\omega}_{\text{space}}$ Hyperparameter combination search space $\chi(\cdot)$ Probability density function for hyperparameter configurations where $y$ is better than the threshold $y^{*}$ $\gamma(\cdot)$ Probability density function for hyperparameter configurations where $y$ is worse than the threshold $y^{*}$ $\hat{\mathbf{X}}^{l,i}_{\text{2D}}$ 2D tensor processed by Inception $\hat{\mathbf{X}}^{l,i}_{\text{1D}}$ 1D sequence after reshaping back $\hat{\textbf{{x}}}^{n}(t)$ The forecasted value $\hat{A}^{l-1}_{f_{1}}$ Weights for sequence aggregation $\textbf{{cD}}_{j}$ Average detail coefficients at the $j$-th level for all sequences $\textbf{{x}}^{n}(t)$ The time point on the $n$-th sequence in $\textbf{{X}}_{\text{1D}}$ $\textbf{{X}}_{\text{1D}}$ Multivariate input time series $\textbf{{X}}_{\text{1D}}^{l}$ Features of the $l$-th layer of $X_{\text{1D}}$ $\textbf{{X}}_{\text{2D}}^{l,i}$ 2D tensor obtained by reshaping $\xi_{j}$ Average amplitude at the $j$-th level $A_{f_{i}}$ Amplitude corresponding to $f_{i}$ $f_{i}$ The $i$-th frequency $k$ Top $k$ amplitudes are selected NN Number of sequences $p_{i}$ The $i$-th period $T$ Number of sampling points per sequence $y$ Loss function BSL Below Sea Level DWT Discrete Wavelet Transform FFT Fast Fourier Transform MAE Mean Absolute Error MAPE Mean Absolute Percentage Error MSE Mean Squared Error MSPE Mean Squared Percentage Error RS Random Sampler TCS Tidal Current Speed TCSF Tidal Current Speed Forecasting TPE Tree-structured Parzen Estimator WECN Wavelet-Enhanced Convolutional Network
@@ 1.2 Literature review
In 1883, Sir G. H. Darwin proposed the harmonic analysis method [10]. He stated, “The tidal oscillations of the ocean can be represented as the sum of several simple harmonic motions.” The harmonic analysis method was later developed by Doodson, who established the famous tidal harmonic component theory. Harmonic models rely on parameter estimation. Doodson continued this research and proposed estimating the parameters of the harmonic model with the least squares method [11]. Earlier technologies did not allow equipment to measure TCS. Research [12] applied harmonic models for TCSF latter. The harmonic model takes into account the multi-periodicity of TCS. However, the harmonic model cannot adapt to the local periodic changes of TCS. There were more models for TCSF as computing power increased.
@@
The emergence of supercomputers brought advancements to physical models. Based on differential equations, physical models use two-dimensional or three-dimensional hydrodynamic models to make forecasts [13]. Physical models can adapt to the non-stationarity of TCS. However, modeling with differential equations depends on physical knowledge. Additionally, solving these equations requires continuous measurement of boundary conditions. Therefore, data-driven models are needed for TCSF.
@@
Statistical models are suitable as a data-driven approach for TCSF. They avoid modeling based on specific TCS measurement locations. However, statistical models rely on strict stationarity assumptions. Therefore, statistical methods have some limitations and they have relatively high requirements in terms of data size, stability of time series, data distribution, and so on [14].
@@
With the rise of artificial intelligence technology, a large number of forecasting methods based on machine learning technology have emerged [15]. Research [16, 17, 18] indicates that machine learning models outperform both statistical and harmonic models in predictive performance. However, machine learning models cannot determine data features. Feature extraction usually requires intuitive knowledge for quantification and expansion.
@@
Deep learning models can provide accurate renewable energy generation forecasts [19]. Deep learning models are applied in the renewable energy field to reduce the complexity of feature engineering. Research [20] used Long Short-Term Memory (LSTM) neural networks for tidal level forecasting. However, deep learning models are still underutilized in the tidal energy field. In other words, while deep learning methods have certain advantages, they are not designed for tidal current time series.
@@ 1.3 Research gaps and proposed solution
Existing research still has limitations. Tidal current time series demonstrate multi-periodicity [21]. These time series include multiple periodic components, covering both astronomical cycles and local periods. Astronomical cycles include diurnal and semi-diurnal tides, among others. In addition to astronomical cycles, there are also local periodicities in tidal currents, manifested as local variations. For example, interactions between tides and waves cause local changes [22]. Multiple overlapping and interacting cycles complicate the model of tidal current time series. Multi-periodicity presents an obstacle to TCSF. Secondly, existing research lacks deep learning models designed for tidal current time series. Therefore, using deep learning methods to consider the multi-periodicity in tidal current time series has sparked our research interest.
We are inspired by the TimesNet model proposed by Wu et al. [23]. The model focuses on multi-periodicity. The model uses the Fast Fourier Transform (FFT) to extract multiple periods from the time series. Based on the periods, the model transforms each 1D sequence into a 2D tensor. Then, the model uses visual backbone networks to capture the 2D variations in the time series. However, the framework lacks consideration for the local periodicity of tidal currents. Research [24] used time series decomposition methods to handle time series. The research inspired us: Discrete Wavelet Transform (DWT) can provide local information on signals in both time and frequency domains. We can handle the series at different scales to capture local periodicity. Additionally, the DWT module helps address local periodicity. The key challenge lies in integrating local periodicity into the learning of multi-periodicity. We propose a framework, Wavelet-Enhanced Convolutional Network (WECN), to address the aforementioned challenges. We use DWT to decompose the time series and measure sub-band amplitudes to identify dominant periods, feeding them into the predictor without reconstructing components. Finally, we consider the forecasting process to be an optimization problem. However, as the dimensionality of hyperparameters increases, the Gaussian model becomes increasingly complex. The Tree-structured Parzen Estimator (TPE) algorithm optimizes hyperparameters to handle the high dimensionality of the hyperparameter space [25].
@@ 2.1 Multi-scale time series analysis
The detail layers with the largest amplitudes show the most significant frequency changes in the signal. Inspired by other research, to avoid meaningless noise from high frequencies, the top $k$ detail layers with the largest amplitudes are selected [26]. $k$ is one of the hyperparameters to be optimized. If $J$ is less than $k$, all amplitudes are selected to form the amplitude array $\boldsymbol{\xi}=[\xi_{1},\xi_{2},\ldots,\xi_{J}]$, and $k$ is updated to $J$. If $J$ is greater than or equal to $k$, the top $k$ values are selected to form the amplitude array $\boldsymbol{\xi}=\text{top}(\xi_{1},\xi_{2},\ldots,\xi_{J})$. Here, $\text{top}(\cdot)$ denotes the top $k$ elements selection to form a new array. Next, the index array $\psi=\text{index}(\xi)$ is obtained, where $\text{index}(\cdot)$ denotes the indices of the elements in the set. Finally, the frequency $f_{i}$ and period $p_{i}$ for the corresponding levels are calculated, expressed as:
@@ 2.2 Tidal current speed predictor
We use WECN to forecast after we obtain the periodic information. Existing methods have limitations in capturing multi-periodicity. Existing methods are either limited by one-dimensional input sequences, making it difficult to capture multi-periodicity, or require complex feature engineering [27]. This research employs the WECN framework to model tidal current time series, as it can fully utilize periodic information for forecasting. The WECN framework uses the TimesNet model proposed by Wu et al. as the infrastructure of the predictor. Wu et al. [23] observed that the time point variation for each period is influenced by the temporal patterns of its neighboring region and changes in its adjacent periods. Therefore, the two-dimensional changes in the time series can be represented as intra-period and inter-period variations.
@@
The adjacent TimesBlock structures are connected using the Residual structure [28]. The WECN framework is illustrated in Figure 4.
@@
Next, we need to merge the $k$ different 1D representations $\{\hat{\mathbf{X}}^{l,1}_{\text{1D}},\cdots,\hat{\mathbf{X}}^{l,k}_{\text{1D}}\}$ for the subsequent layer. The amplitude $A$ reflects the relative importance of the selected frequencies and periods, which correlates with the significance of each transformed 2D tensor [30]. Therefore, the Softmax function calculates the weights. After this, the 1D representations based on amplitude are aggregated.
@@
$\textbf{{X}}_{\text{1D}}$ is processed through each TimesBlock with residual connections. Then, the features from the final layer are projected into the output space via a fully connected layer for multi-step forecasting. Single-step forecasting is sometimes insufficient to ensure the reliability and controllability of tidal energy systems. Therefore, this paper adopts and discusses the multi-step forecasting mechanism, providing more information for future decision-making [31]. We use a recursive forecasting strategy for multi-step forecasting, where the current output forecasting serves as the input for the next step. The multi-step forecasting process can be seen in Figure 5. The multi-step forecast results belong to $\mathbb{R}^{T\times N\times\beta}$, where $\beta$ denotes the output length. The $g$-th step forecasting value for the $t$-th sampling point in the $n$-th sequence can be represented as $\hat{x}^{n}(t+g)$, where $g=1,2,...,\beta$.
@@ 2.3 The process of hyperparameter optimization
We recognize the importance of selecting optimal hyperparameters. To achieve this, we use the Tree-structured Parzen Estimator (TPE). TPE is a variant of Bayesian optimization suitable for high-dimensional and discrete parameter spaces. Unlike traditional Bayesian methods that rely on Gaussian processes, TPE improves efficiency by constructing probability density functions in the hyperparameter search space using Parzen window density estimation [32]. Next, we transform the forecasting problem into an optimization process, with the design of the objective function discussed in Section 2.3.1 and the application of TPE in Section 2.3.2.

### v4

@@ Abstract
Tidal energy is one of the key components in increasing the penetration of renewable energy. High tidal energy penetration into the electrical grid depends on accurate tidal current speed forecasting. Model inaccuracies hinder forecast accuracy. Previous research primarily used physical models to forecast tidal current speed, yet tidal current variations influenced by the orbital periods of celestial bodies make accurate physical modeling challenging. Research on the multi-periodicity of tides is crucial for forecasting tidal current speed. We propose the Wavelet-Enhanced Convolutional Network to learn multi-periodicity. The framework embeds intra-period and inter-period variations of one-dimensional tidal current data into the rows and columns, respectively, of a two-dimensional tensor. Then, the two-dimensional variations of the sequence can be processed by convolutional kernels. We integrate a time-frequency analysis method into the framework to further address local periodic features. Additionally, to enhance the framework’s stability, we optimize the framework’s hyperparameters with the Tree-structured Parzen Estimator. The proposed framework captures multi-periodic dependencies in tidal current data. Numerical results show a 10-step average Mean Absolute Error of 0.025, with at least a 1.44% error reduction compared to other baselines. Further ablation studies show a 1.4% reduction in Mean Absolute Percentage Error on the data with artificially added periodic fluctuations.
@@ 1.1 Background
The development of renewable energy can promote emission reduction [CAGLAYAN2019113794]. Tidal energy is considered one of the most competitive forms of renewable energy [OROURKE2014726].
Tidal energy attracts the attention of many coastal countries [SHAO2023125476], especially the UK. The potential of tidal energy lies in three aspects: cost-effectiveness, stability, and predictability. As tidal energy gains more attention, its cost-effectiveness is expected to improve, as shown in Figure 1. Tidal energy may replace other renewable energies when the Levelized Cost of Energy (LCOE) falls below £49-55/MWh (expected by 2050) [tiger2022cost]. Additionally, tidal energy offers high stability, which is crucial for the economic feasibility of any renewable energy project [GRABBE20091898]. Therefore, the introduction of tidal energy can enhance grid stability. Moreover, tidal energy provides the advantage of predictability over large time scales. Consequently, tidal energy is an excellent choice as a base-load energy supply [OROURKE2014726].
Tidal current turbines generate electricity using the kinetic energy of tidal currents, like wind turbines use the kinetic energy of air [YANG2024119603]. Therefore, fluctuations in Tidal Current Speed (TCS) reduce grid stability [MONAHAN2023103596]. Enhancing the accuracy of short-term forecasting helps mitigate the impact of TCS fluctuations on grid stability, thereby increasing the penetration of tidal energy into the grid [6665108]. Research [wang2023wavelet] proposes to explore the statistical properties of real-life time series and design customized deep learning models for high-performance time series forecasting. The research on statistical properties of time series inspires us to explore the multi-periodicity within the TCS series. We increase the consideration of the statistical properties of TCS to improve the accuracy of TCSF.
@@ 1.2 Literature review
In 1883, Sir G. H. Darwin proposed the harmonic analysis method [Darwin_2009]. He stated, “The tidal oscillations of the ocean can be represented as the sum of several simple harmonic motions.” The harmonic analysis method was later developed by Doodson, who established the famous tidal harmonic component theory. Harmonic models rely on parameter estimation. Doodson continued this research and proposed estimating the parameters of the harmonic model with the least squares method [doodson1957analysis]. Earlier technologies did not allow equipment to measure TCS. Research [Fornerino_LeProvost_2015] applied harmonic models for TCSF latter. The harmonic model takes into account the multi-periodicity of TCS. However, the harmonic model cannot adapt to the local periodic changes of TCS. There were more models for TCSF as computing power increased.
@@
The emergence of supercomputers brought advancements to physical models. Based on differential equations, physical models use two-dimensional or three-dimensional hydrodynamic models to make forecasts [owen1980three]. Physical models can adapt to the non-stationarity of TCS. However, modeling with differential equations depends on physical knowledge. Additionally, solving these equations requires continuous measurement of boundary conditions. Therefore, data-driven models are needed for TCSF.
@@
Statistical models are suitable as a data-driven approach for TCSF. They avoid modeling based on specific TCS measurement locations. However, statistical models rely on strict stationarity assumptions. Therefore, statistical methods have some limitations and they have relatively high requirements in terms of data size, stability of time series, data distribution, and so on [MA2020120159].
@@
With the rise of artificial intelligence technology, a large number of forecasting methods based on machine learning technology have emerged [YANG2022119849]. Research [5589321, sarkar2016machine, sarkar2018prediction] indicates that machine learning models outperform both statistical and harmonic models in predictive performance. However, machine learning models cannot determine data features. Feature extraction usually requires intuitive knowledge for quantification and expansion.
@@
Deep learning models can provide accurate renewable energy generation forecasts [Li2024]. Deep learning models are applied in the renewable energy field to reduce the complexity of feature engineering. Research [9169644] used Long Short-Term Memory (LSTM) neural networks for tidal level forecasting. However, deep learning models are still underutilized in the tidal energy field. In other words, while deep learning methods have certain advantages, they are not designed for tidal current time series.
@@ 1.3 Research gaps and proposed solution
Existing research still has limitations. Tidal current time series demonstrate multi-periodicity [qian2022tidal]. These time series include multiple periodic components, covering both astronomical cycles and local periods. Astronomical cycles include diurnal and semi-diurnal tides, among others. In addition to astronomical cycles, there are also local periodicities in tidal currents, manifested as local variations. For example, interactions between tides and waves cause local changes [jay1999comparison]. Multiple overlapping and interacting cycles complicate the model of tidal current time series. Multi-periodicity presents an obstacle to TCSF. Secondly, existing research lacks deep learning models designed for tidal current time series. Therefore, using deep learning methods to consider the multi-periodicity in tidal current time series has sparked our research interest.
We are inspired by the TimesNet model proposed by Wu et al. [wu2022timesnet]. The model focuses on multi-periodicity. The model uses the Fast Fourier Transform (FFT) to extract multiple periods from the time series. Based on the periods, the model transforms each 1D sequence into a 2D tensor. Then, the model uses visual backbone networks to capture the 2D variations in the time series. However, the framework lacks consideration for the local periodicity of tidal currents. Research [ZHANG2022611] used time series decomposition methods to handle time series. The research inspired us: Discrete Wavelet Transform (DWT) can provide local information on signals in both time and frequency domains. We can handle the series at different scales to capture local periodicity. Additionally, the DWT module helps address local periodicity. The key challenge lies in integrating local periodicity into the learning of multi-periodicity. We propose a framework, Wavelet-Enhanced Convolutional Network (WECN), to address the aforementioned challenges. We use DWT to decompose the time series and measure sub-band amplitudes to identify dominant periods, feeding them into the predictor without reconstructing components. Finally, we consider the forecasting process to be an optimization problem. However, as the dimensionality of hyperparameters increases, the Gaussian model becomes increasingly complex. The Tree-structured Parzen Estimator (TPE) algorithm optimizes hyperparameters to handle the high dimensionality of the hyperparameter space [KIM2024118010].
@@ 2.1 Multi-scale time series analysis
The detail layers with the largest amplitudes show the most significant frequency changes in the signal. Inspired by other research, to avoid meaningless noise from high frequencies, the top $k$ detail layers with the largest amplitudes are selected [zhou2022fedformer]. $k$ is one of the hyperparameters to be optimized. If $J$ is less than $k$, all amplitudes are selected to form the amplitude array $\boldsymbol{\xi}=[\xi_{1},\xi_{2},\ldots,\xi_{J}]$, and $k$ is updated to $J$. If $J$ is greater than or equal to $k$, the top $k$ values are selected to form the amplitude array $\boldsymbol{\xi}=\text{top}(\xi_{1},\xi_{2},\ldots,\xi_{J})$. Here, $\text{top}(\cdot)$ denotes the top $k$ elements selection to form a new array. Next, the index array $\psi=\text{index}(\xi)$ is obtained, where $\text{index}(\cdot)$ denotes the indices of the elements in the set. Finally, the frequency $f_{i}$ and period $p_{i}$ for the corresponding levels are calculated, expressed as:
@@ 2.2 Tidal current speed predictor
We use WECN to forecast after we obtain the periodic information. Existing methods have limitations in capturing multi-periodicity. Existing methods are either limited by one-dimensional input sequences, making it difficult to capture multi-periodicity, or require complex feature engineering [jia2024witran]. This research employs the WECN framework to model tidal current time series, as it can fully utilize periodic information for forecasting. The WECN framework uses the TimesNet model proposed by Wu et al. as the infrastructure of the predictor. Wu et al. [wu2022timesnet] observed that the time point variation for each period is influenced by the temporal patterns of its neighboring region and changes in its adjacent periods. Therefore, the two-dimensional changes in the time series can be represented as intra-period and inter-period variations.
@@
The adjacent TimesBlock structures are connected using the Residual structure [he2016deep]. The WECN framework is illustrated in Figure 4.
@@
Next, we need to merge the $k$ different 1D representations $\{\hat{\mathbf{X}}^{l,1}_{\text{1D}},\cdots,\hat{\mathbf{X}}^{l,k}_{\text{1D}}\}$ for the subsequent layer. The amplitude $A$ reflects the relative importance of the selected frequencies and periods, which correlates with the significance of each transformed 2D tensor [NEURIPS2021_bcc0d400]. Therefore, the Softmax function calculates the weights. After this, the 1D representations based on amplitude are aggregated.
@@
$\textbf{{X}}_{\text{1D}}$ is processed through each TimesBlock with residual connections. Then, the features from the final layer are projected into the output space via a fully connected layer for multi-step forecasting. Single-step forecasting is sometimes insufficient to ensure the reliability and controllability of tidal energy systems. Therefore, this paper adopts and discusses the multi-step forecasting mechanism, providing more information for future decision-making [du2019novel]. In practical tidal energy systems, forecasts are typically updated in an online and rolling manner as new observations become available. Recursive forecasting enables low-latency step-by-step prediction while supporting flexible extension of the forecasting horizon. Under such real-time operational constraints, we use a recursive forecasting strategy for multi-step forecasting, where the current output forecasting serves as the input for the next step. The multi-step forecasting process can be seen in Figure 5. The multi-step forecast results belong to $\mathbb{R}^{T\times N\times\beta}$, where $\beta$ denotes the output length. The $g$-th step forecasting value for the $t$-th sampling point in the $n$-th sequence can be represented as $\hat{x}^{n}(t+g)$, where $g=1,2,...,\beta$.
2.3 Interpretation of the tidal receptive field
Tidal current speed contains (i) phase evolution within one tidal cycle and (ii) cycle-to-cycle modulation across consecutive cycles. Equation (7) reshapes $\textbf{{X}}_{\text{1D}}^{l-1}$ into $\textbf{{X}}_{\text{2D}}^{l,i}$ so that the 2D receptive field used in Equation (9) aligns with these two physical structures: the row axis corresponds to intra-period (within-cycle) phase neighborhoods, and the column axis corresponds to inter-period (across-cycle) neighborhoods.
We define an index mapping from the 2D coordinates of $\textbf{{X}}_{\text{2D}}^{l,i}$ to the original 1D position:
$\tau(r,c)=(c-1)p_{i}+r,\qquad r\in\{1,\ldots,p_{i}\},\ c\in\{1,\ldots,q_{i}\}.$ (11)
Thus, moving along $r$ captures local phase-shape changes within one cycle, while moving along $c$ compares similar phases across neighboring cycles.
If a kernel in Equation (9) covers an $a\times b$ patch on $\textbf{{X}}_{\text{2D}}^{l,i}$, then the corresponding maximum span on the original 1D index is
$\Delta\tau_{\max}=(b-1)p_{i}+(a-1),$ (12)
and the physical time span equals $\Delta\tau_{\max}\Delta_{s}$, where $\Delta_{s}$ is the sampling interval. Therefore, $a$ controls the within-cycle (phase-neighborhood) coverage and $b$ controls the across-cycle coverage for capturing cycle-to-cycle modulation and slow drift.
@@ 2.4 The process of hyperparameter optimization
We recognize the importance of selecting optimal hyperparameters. To achieve this, we use the Tree-structured Parzen Estimator (TPE). TPE is a variant of Bayesian optimization suitable for high-dimensional and discrete parameter spaces. Unlike traditional Bayesian methods that rely on Gaussian processes, TPE improves efficiency by constructing probability density functions in the hyperparameter search space using Parzen window density estimation [ALISALAMAI2023118658]. Next, we transform the forecasting problem into an optimization process, with the design of the objective function discussed in Section 2.4.1 and the application of TPE in Section 2.4.2.
@@ 3.1 Experimental setup
Time series forecasting is highly susceptible to look-ahead bias, particularly when future information leaks into preprocessing, model training, or model selection. In this work, we strictly follow a modeling protocol:
  • The training, validation, and testing sets are split in chronological order, ensuring that the validation and test segments occur strictly after the training segment;
  • No future samples are used during preprocessing, i.e., any parameters or statistics required by preprocessing are estimated using the training segment only and then applied unchanged to the validation and test sets;
  • Training samples are constructed using causal input windows, such that each target value is predicted solely from historical observations.
This protocol ensures that the reported performance faithfully reflects realistic forward-in-time forecasting.